blakea0801:

The point P(21, 28) is on the terminal side of θ. Evaluate sin θ.

Gdeinward:

Hello, WELCOME TO QUESTION COVE!!!!! I cannot help you on this question. However, I recommend @dude @Vocaloid @TheSmartOne

dude:

This is fairly straightforward  Sin = \(\large\frac{opposite}{hypotenuse}\) We just need to find the hypotenuse We know the side lengths based on the point location, so use Pythagorean Theorem to solve for c c=hypotenuse \(21^2+28^2=c^2\)